# Multipartite Entanglement Distribution in Quantum Networks using Subgraph Complementations

### v4

@@ VII-B Complete bipartite graph state generation
Figure 7: Comparison of fidelity of (a) 6-qubit complete bipartite graph state and (b) 20-qubit complete bipartite graph state, distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.003$, with the results obtained over 10000 trials for each of the graph states. The red line indicates a fidelity of 0.75.
We expect the difference in fidelity to monotonically increase in favor of the subgraph complementation algorithm as the number of qubits increase. We conjecture this difference is a consequence of the $O(n)$ scaling in time and gate operations for our algorithm as compared to $O(n^{2})$ for [8]. For a smaller number of qubits, a linear overhead due to the edge reset operation results in a higher number of operations to carry out a subgraph complementation system as compared to locally creating a graph state and teleporting all the qubits. However, as the number of qubits increases asymptotically, the fidelity to distribute a complete bipartite graph state for our algorithm should increase only linearly as compared to quadratically for [8]. We observe that Fig. 8 exhibits this trend.
Figure 8: Comparison of fidelity of an n-qubit complete bipartite graph state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] varying $6\leq n\leq 30$ with $p=10^{-3}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
VII-C Parallelized GHZ state generation
We also carried out simulations of distributing a GHZ state with parallelized gate operations as described in section VI. We compare our parallelized SC operation protocol to a parallelized version of the algorithm by Avis et al., in which a complete graph is created locally in the central node by parallelizing $CZ$ gates according to an edge coloring of the complete graph [23]. This algorithm requires $O(n)$ time steps to distribute an $n$-qubit GHZ state, however the parallelized subgraph complementation protocol only requires $O(\log n)$ time steps. Fig. 9(a) compares the fidelity produced by our parallelized SC algorithm with that produced by the parallelized factory node + teleportation algorithm for a 6-qubit GHZ state, as a function of error probability. The parallelized SC algorithm uses 2 auxiliary qubits.
@@ VII-C Parallelized GHZ state generation
Figure 9: Comparison of fidelity of (a) 6-qubit GHZ state and (b) 20-qubit GHZ state, distributed according to the parallelized subgraph complementation algorithm (with (a) 2 auxiliary qubits and (b) 4 auxiliary qubits respectively) and the factory node + teleportation algorithm having $0\leq p\leq 0.015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
@@
Figure 10: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the parallelized subgraph complementation algorithm and the factory node + teleportation algorithm varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
It is interesting to note that on comparing Fig. 6 and Fig. 10, we observe that the fidelity obtained from distributing $n$-qubit GHZ states without parallelization is higher than with parallelization (although both outperform the parallelized factory node + teleportation algorithm). This is expected since we are using $n/\log(n)$ extra qubits for the parallelized protocol. However, we are also gaining an exponential decrease in distribution time ($O(\log n)$ with parallelization as compared to $O(n)$ without it). Therefore, this tradeoff would be favourable in most cases, especially when considering memory decoherence of the qubits which would increase proportionally with distribution time.
VIII Conclusion
We have presented a novel and efficient protocol for distributing an arbitrary graph state across a quantum network by drawing upon the mathematical theory of subgraph complementations, which we show to be strongly linked to the entanglement and structure of the graph state. We show, through simulations, that this protocol generates graph states having high fidelity in the presence of noisy operations. This algorithm also provides a natural framework to create a graph state in a distributed fashion, since subgraph complementations commute and it may be possible to implement several $SC$ operations in parallel (since their order does not matter) and merge the subgraphs at the end.
This could also be extended to a procedure for distributing arbitrary stabilizer states, since graph states are equivalent to stabilizer states up to local operations. We anticipate that the framework of subgraph complementations will be useful for other applications as well, for instance quantum error correction and secret sharing. Further work includes incorporating error correction in estimating runtime and fidelity through noisy simulations of our algorithm, considering the detailed impact of memory decoherence, as well as optimizing our protocol for more domain-specific purposes.
@@ Appendix A Additional simulation results
Figure 11: Comparison of fidelity of a 5-qubit (top) and 20-qubit (bottom) GHZ state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.0015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 12: (Simulation considers noisy gate operations and memory decoherence) Comparison of fidelity of a 3-qubit GHZ state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] having $0\leq p\leq 0.0015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 13: (Simulation considers noisy gate operations and memory decoherence) Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the subgraph complementation algorithm and the factory node + teleportation algorithm [8] varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.

### v5

@@ 7.2 Complete bipartite graph state generation
Figure 7: Comparison of fidelity of (a) 6-qubit complete bipartite graph state and (b) 20-qubit complete bipartite graph state, distributed according to the SC algorithm (4.1) and the FNT algorithm [2] having $0\leq p\leq 0.003$, with the results obtained over 10000 trials for each of the graph states. The red line indicates a fidelity of 0.75.
We expect the difference in fidelity to monotonically increase in favor of the SC algorithm as the number of qubits increase. We conjecture this difference is a consequence of the $O(n)$ scaling in time and gate operations, or circuit depth, for our algorithm as compared to $O(n^{2})$ for FNT (1.1). For a smaller number of qubits, a linear overhead due to the edge reset operation results in a higher number of operations to carry out a subgraph complementation system as compared to locally creating a graph state and teleporting all the qubits. However, as the number of qubits increases asymptotically, the fidelity to distribute a complete bipartite graph state for our algorithm should increase only linearly as compared to quadratically for FNT (1.1). We observe that Fig. 8 exhibits this trend with the crossover in fidelities observed at $n=12$. Note that we set $p=10^{-3}$ in the figure as compared to $p=10^{-4}$ in Fig. 6 since at $p=10^{-4}$ complete bipartite graph states (as opposed to GHZ states) created by both algorithms are less noisy and too close in fidelity for a meaningful comparison. This is our guiding principle for choosing parameters in all figures, with comparisons expected to generalize in other parameter regimes.
Figure 8: Comparison of fidelity of an $n$-qubit complete bipartite graph state distributed according to the SC algorithm (4.1) and the FNT algorithm [2] varying $6\leq n\leq 30$ with $p=10^{-3}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
7.3 Parallelized GHZ state generation
We also carried out simulations of distributing a GHZ state with parallelized gate operations as described in section VI. We compare our parallelized SC algorithm (6.2) to a parallelized version of the FNT algorithm (1.1), in which a complete graph is created locally in the central node by parallelizing $CZ$ gates according to an edge coloring of the complete graph [7]. This algorithm requires $O(n)$ time steps to distribute an $n$-qubit GHZ state, however the parallelized subgraph complementation algorithm only requires $O(\log n)$ time steps. Fig. 8(a) compares the fidelity produced by our parallelized SC algorithm with that produced by the parallelized FNT algorithm for a 6-qubit GHZ state, as a function of error probability. The parallelized SC algorithm uses 2 auxiliary qubits.
@@ 7.3 Parallelized GHZ state generation
Figure 9: Comparison of fidelity of (a) 6-qubit GHZ state and (b) 20-qubit GHZ state, distributed according to the parallelized SC algorithm (with (a) 2 auxiliary qubits and (b) 4 auxiliary qubits respectively) and the FNT algorithm having $0\leq p\leq 0.015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
@@
Figure 10: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the parallelized SC algorithm and the FNT algorithm varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
It is interesting to note that on comparing Fig. 6 and Fig. 10, we observe that the fidelity obtained from distributing $n$-qubit GHZ states without parallelization is higher than with parallelization (although both outperform the parallelized FNT algorithm). This is expected since we are using $n/\log(n)$ extra qubits for the parallelized algorithm. However, we are also gaining an exponential decrease in distribution time ($O(\log n)$ with parallelization as compared to $O(n)$ without it). Therefore, this tradeoff would be favourable in most cases, especially when considering memory decoherence of the qubits which would increase proportionally with distribution time.
8 Conclusion
We have presented a novel and efficient algorithm for distributing an arbitrary graph state across a quantum network by drawing upon the mathematical theory of subgraph complementations, which we show to be strongly linked to the entanglement and structure of the graph state. We show, through simulations, that this algorithm generates graph states having high fidelity in the presence of noisy operations. This algorithm also provides a natural framework to create a graph state in a distributed fashion, since subgraph complementations commute and it may be possible to implement several $SC$ operations in parallel (since their order does not matter) and merge the subgraphs at the end.
This could also be extended to a procedure for distributing arbitrary stabilizer states, since graph states are equivalent to stabilizer states up to local operations. We anticipate that the framework of subgraph complementations will be useful for other applications as well, for instance quantum error correction and secret sharing. Further work includes incorporating error correction in estimating runtime and fidelity through noisy simulations of our algorithm, considering the detailed impact of memory decoherence, as well as optimizing our algorithm for more domain-specific purposes.
9 Acknowledgements
This work was funded by the Army Research Office (ARO) MURI on Quantum Network Science under grant number W911NF2110325, NSF-ERC Center for Quantum Networks grant EEC-194158 and NSF under grant 2402861.
@@ Appendix A Additional simulation results
Figure 11: Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the SC algorithm (4.1), the FNT algorithm (Bipartite A protocol) (1.1), and the Bipartite B protocol, considering (top) and neglecting (bottom) noise from local correction unitaries, varying $6\leq n\leq 30$ with $p=10^{-3}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.
In this appendix, we provide some additional simulation results. Figure 11 shows the fidelity of $n$-qubit GHZ graph states generated using three different algorithms: the SC algorithm, the FNT algorithm (Bipartite A), and the Bipartite B protocol, for $6\leq n\leq 30$ with error probability $p=10^{-3}$. The top plot includes noise from local correction unitaries, while the bottom plot neglects this noise. We observe that the SC algorithm consistently achieves higher fidelity than the other algorithms across all $n$ values, even when local correction noise is accounted for. While the Bipartite B protocol requires fewer gate operations, the Bipartite A protocol outperforms it in fidelity, diverging from earlier results reported by Cuquet and Calsamiglia [10]. Interestingly, when local correction noise is excluded (bottom plot), we recover those earlier trends, showing improved performance of Bipartite B, consistent with assumptions made in prior work. These findings highlight the significance of accounting for realistic noise sources and support the robustness of the SC algorithm under more comprehensive noise models.
Figure 12: Comparison of fidelity of a 5-qubit (top) and 20-qubit (bottom) GHZ state distributed according to the SC algorithm (4.1) and the FNT algorithm [2] having $0\leq p\leq 0.0015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figure 12 compares the fidelity of GHZ states of 5 qubits (top) and 20 qubits (bottom) generated using the SC and FNT algorithms as a function of the noise parameter $0\leq p\leq 0.0015$. In both cases, the SC algorithm outperforms the FNT algorithm across the entire range of noise values. Notably, the gap in fidelity between the two algorithms becomes more pronounced in the 20-qubit scenario.
Figure 13: (Simulation considers noisy gate operations and memory decoherence) Comparison of fidelity of a 3-qubit GHZ state distributed according to the SC algorithm (4.1) and the FNT algorithm [2] having $0\leq p\leq 0.0015$, with the results obtained over 10000 trials. The red line indicates a fidelity of 0.75.
Figures 13 and 14 present the fidelity of GHZ states distributed using the SC and FNT algorithms under a more realistic noise model that includes both gate errors and memory decoherence. In Figure 13, we consider a 3-qubit GHZ state with varying noise levels $0\leq p\leq 0.0015$, while in Figure 14, we examine $n$-qubit GHZ graph states for $6\leq n\leq 30$ with a fixed error rate of $p=10^{-4}$. In both scenarios, the SC algorithm consistently outperforms the FNT algorithm, though the difference does not seem to increase significantly with increasing number of qubits, unlike in other cases which did not consider memory decoherence. We conjecture this might be a consequence of a larger scaling being required with respect to the number of qubits in this case, since the noise dominates with a smaller number of qubits.
Figure 14: (Simulation considers noisy gate operations and memory decoherence) Comparison of fidelity of an $n$-qubit GHZ graph state distributed according to the SC algorithm (4.1) and the FNT algorithm [2] varying $6\leq n\leq 30$ with $p=10^{-4}$, with the results obtained over 5000 trials. The red line indicates a fidelity of 0.75.